A chemist needs 130 milliliters of a 39% solution but has only 3% and 55% solutions available. How many milliliters of each should be mixed to get the desired solution?
Answer by mananth(16946)

---------percent ---------------- quantity
Solution I 3 ---------------- x ml
Solution II 55 ------ 130 - x ml
Mixture 39.00% ---------------- 130

3 x + 55 ( 130 - x ) = 39 * 130

3x+7150 -55x=5070
3x-55x=5070-7150
-52x= -2080
/ -52
x=40 ml 3.00% Solution I
90 ml 55.00% Solution II